### v1

Abstract
Hyperspectral image (HSI) fusion addresses the challenge of reconstructing High-Resolution HSIs (HR-HSIs) from High-Resolution Multispectral images (HR-MSIs) and Low-Resolution HSIs (LR-HSIs), a critical task given the high costs and hardware limitations associated with acquiring high-quality HSIs. While existing methods leverage spatial and spectral relationships, they often suffer from limited receptive fields and insufficient feature utilization, leading to suboptimal performance. Furthermore, the scarcity of high-quality HSI data highlights the importance of efficient data utilization to maximize reconstruction quality. To address these issues, we propose HyFusion, a novel framework designed to enhance the receptive field and enable effective feature map reusing, thereby maximizing data utilization. First, HR-MSI and LR-HSI inputs are concatenated to form a quasi-fused draft, preserving complementary spatial and spectral details. Next, the Enhanced Reception Field Block (ERFB) is introduced, combining shifting-window attention and dense connections to expand the receptive field, effectively capturing long-range dependencies and reusing features to reduce information loss, thereby boosting data efficiency. Finally, the Dual-Coupled Network (DCN) dynamically extracts high-frequency spectral and spatial features from LR-HSI and HR-MSI, ensuring efficient cross-domain fusion. Extensive experiments demonstrate that HyFusion achieves state-of-the-art performance in HR-MSI/LR-HSI fusion, significantly improving reconstruction quality while maintaining a compact model size and computational efficiency. By integrating enhanced receptive fields and feature map reusing, HyFusion provides a practical and effective solution for HSI fusion in resource-constrained scenarios, setting a new benchmark in hyperspectral imaging. Our code will be publicly available.
Index Terms:
Hyperspectral Image Fusion, Hyperspectral Image Pansharping, Transformer, Reception Field, Data Efficiency
I Introduction
Hyperspectral imaging (HSI) captures detailed spectral information for every pixel in an image, providing a significantly broader spectral range compared to conventional imaging techniques, with channels ranging from tens to hundreds. This unique capability has enabled diverse applications across fields such as science, military, agriculture, and medicine [1]. However, HSI sensing systems face notable hardware constraints, particularly in miniaturized satellites and resource-limited platforms. These constraints limit the achievable spectral and spatial resolutions. As a result, capturing high-quality, high-resolution hyperspectral images (HR-HSI) is both technically challenging and financially costly, particularly for resource-constrained platforms such as small satellites. Consequently, data scarcity poses a significant hurdle for the practical deployment of many HSI applications.
One widely adopted solution to address these limitations is to capture high-resolution multispectral images (HR-MSIs) alongside low-resolution hyperspectral images (LR-HSIs). Fusion-based super-resolution (SR) techniques [2, 3, 4] combine HR-MSIs and LR-HSIs to reconstruct high-resolution hyperspectral images (HR-HSIs). These techniques have proven effective in overcoming resolution limitations and maximizing the utility of limited HSI data [5, 6, 7, 8, 9].
Despite their effectiveness, HSI fusion tasks remain inherently challenging, as they require the seamless integration of spatial and spectral information while preserving fine spectral details and enhancing spatial resolution. Existing methods often face two major challenges. First, the loss of critical spatial and spectral information as network depth increases. Second, the underutilization of rich spatial-spectral relationships in the input data. This information loss is particularly problematic in HSI applications, where making the most of the limited available data is crucial due to the scarcity of high-quality HSI datasets. While deep learning (DL)-based HSI fusion methods [10, 11, 12, 13] have advanced the field, many state-of-the-art techniques fail to fully address key challenges such as receptive field expansion and feature map reuse. Consequently, these methods often fall short in maximizing data utilization, which compromises their overall fusion performance. Furthermore, the computational and hardware limitations inherent to HSI applications demand solutions that are not only highly effective but also efficient and easy to implement.
To overcome these challenges, we propose HyFusion, a novel framework designed to enhance the receptive field and reuse feature maps, enabling more efficient data utilization and improved HR-HSI reconstruction. The receptive field plays a crucial role by capturing long-range dependencies and spatial-spectral correlations, as illustrated in Figure 1. While feature map reuse, facilitated by dense connections, retains and amplifies extracted information throughout the network. At the core of HyFusion is the Enhanced Reception Field Block (ERFB), which integrates Improved Swin Transformer Layers (ISTL) with dense connections to expand the receptive field and preserve critical spatial-spectral features, addressing the limitations of existing methods by mitigating information loss and enhancing feature retention. Additionally, HyFusion introduces a Dual-Coupled Network (DCN) comprising two specialized networks: SpeNet, for extracting spectral features, and SpaNet, for extracting spatial features. To guide network convergence, we employ task-specific spectral-aware and frequency-aware loss functions, which align reconstructed HR-HSIs with ground-truth in both frequency and spatial-spectral domains, ensuring enhanced fidelity and minimizing artifacts in the final reconstructions.
In summary, our key contributions are as follows:
  • We propose the HyFusion framework, which emphasizes Enhanced Reception Field and Feature Map Reusing through the proposed ERFB, enabling the network to maximize data utilization by capturing long-range dependencies and reusing extracted features across layers.
  • The proposed DCN dynamically integrates spectral and spatial features from HR-MSI and LR-HSI, leveraging dense connections to fully exploit limited data while maintaining a simple and efficient network architecture.
  • We design task-specific loss functions to ensure alignment between reconstructed HR-HSIs and ground truth in both frequency and spatial-spectral domains, improving fidelity and eliminating artifacts.
By integrating enhanced receptive fields and feature map reusing into the fusion process, HyFusion effectively addresses the challenges of data scarcity and information loss, establishing a new benchmark for efficient and high-quality HSI fusion in resource-constrained environments.
Figure 1: Visualization of Effective Reception Field. (a) naive Swin Transformer Layer (STL), (RSTB in deep feature extraction [14]); (b) with dense connection (using RDG [15]); (c) with dense connection and improved STL (Ours).
II Proposed Method
Figure 2: The proposed HyFusion network architecture integrates an Enhanced Reception Field Block (ERFB) with an Improved Swin Transformer Layer (ISTL) for HSI/MSI fusion. Through our Dual-Coupled Network (DCN), the architecture learns spatial-spectral representations across two specialized branches. The combination of ERFB, ISTL and DCN enables effective capture of long-range dependencies while maintaining streamlined design, leading to superior fusion results through efficient information exchange between spatial and spectral domains, thereby improving data efficiency for HSI real-world applications.
II-A Architecture of Proposed HyFusion
To design an efficient LR-HSI/HR-MSI fusion model, we aim to use dense-connections [16] to minimize the loss of valuable spectral and spatial information during fusion and enhance the receptive field, thereby ensuring effective information complementarity and high-quality HR-HSI reconstruction. However, cross-domain information fusion requires effective interaction while considering the high computational and hardware resource consumption due to the large size of HSI. Thus, the network architecture must be simple, efficient, and effective for practical applications.
Inspired by ensemble learning in traditional machine learning, which enhances classification performance by using two simple independent classifiers for different modalities and combining their results through voting or weighted averaging, our DCN leverages these advantages. Specifically, we fully exploit the input’s spectral and spatial information by designing two distinct subnetworks to aggregate the information, thereby improving fusion performance with lower computational and spatial complexity.
This section illustrates the architectural design of our DCN, as shown in Figure 2, including the spatial and spectral perception networks (SpaNet and SpeNet) for LR-HSI and HR-MSI. To effectively combine the features of LR-HSI and HR-MSI, the proposed SpaNet and SpeNet not only extract their corresponding feature representations but also utilize early image fusion to allow the corresponding counterparts to provide guidance simultaneously. This approach allows us to seamlessly integrate spatial and spectral feature representations without increasing model complexity.
II-B Proposed Dual-Coupled Network
Let the LR-HSI and HR-MSI be $\mathbf{X}_{h}\in\mathbb{R}^{h_{h}\times w_{h}\times b}$ and $\mathbf{X}_{m}\in\mathbb{R}^{h\times w\times b_{m}}$. The observable LR-HSI can be modeled as $\mathbf{X}_{h}=\mathbf{Y}\mathbf{B}$, where $\mathbf{B}$ is a blurring matrix that reduces the pixel count. Similarly, the observable HR-MSI is given by $\mathbf{X}_{m}=\mathbf{D}\mathbf{Y}$, where $\mathbf{D}$ is a downsampling matrix that decreases the number of spectral bands.
The reconstructed HR-HSI denotes $\mathbf{Y}^{*}\in\mathbb{R}^{w\times h\times b}$ by:
${\mathbf{Z}_{hm}}=f_{\text{DCN}}(\mathbf{X}_{h},\mathbf{X}_{m})=\text{Concat}.(\mathbf{Z}_{h},\mathbf{Z}_{m}),$ (1)
where the high-spectral-resolution and high-spatial-resolution features are obtained as:
$\mathbf{Z}_{h}=f_{\text{SpeNet}}(\mathbf{X}_{h},\mathbf{X}_{m});\\ \mathbf{Z}_{m}=f_{\text{SpaNet}}(\mathbf{X}_{m},\mathbf{X}_{h})$ (2)
For SpeNet, we up-sample the LR-HSI to obtain $\mathbf{X}_{h}^{u}=f_{\text{up}}(\mathbf{X}_{h})$, where $f_{\text{up}}$ denotes the bilinear interpolation function. Then, the quasi-spectral draft is obtained by:
$\mathbf{Z}^{\prime}_{h}=\text{Concat}.(\mathbf{X}_{h}^{u},\mathbf{X}_{m})\vspace{-0.2cm}$ (3)
Similarly for SpaNet, we down-sample the HR-MSI to obtain $\mathbf{X}_{m}^{d}=f_{\text{down}}(\mathbf{X}_{m})$, where $f_{\text{down}}$ denotes the bicubic interpolation function, creating a quasi-spatial draft:
$\mathbf{Z}_{m}^{{}^{\prime}}=\text{Concat}.(\mathbf{X}_{h},\mathbf{X}_{m}^{d})\vspace{-0.2cm}$ (4)
After obtaining these preliminary drafts, shallow feature extraction is performed using $3\times 3$ convolution with LeakyReLU activation (negative slope $0.2$). The ERFB is then used for deep feature extraction, with skip-connections enhancing feature quality and stabilizing training. The reconstructed HR-HSI is:
${\mathbf{Y}^{*}}=f_{\text{Rec}}({\mathbf{Z}_{hm}})\vspace{-0.2cm}$ (5)
TABLE I: Performance evaluation and complexity comparison of the proposed method and other fusion models. The best results are highlighted in bold-red, while the second-best results are highlighted in bold-blue. Rows are colored to distinguish different approaches: CNN-based, Transformer-based, and the proposed methods.
Method Complexity 4 Bands LR-HSI 6 Bands LR-HSI
#Params↓ FLOPs↓ Run-time↓ Memory↓ PSNR↑ SAM↓ RMSE↓ ERGAS↓ PSNR↑ SAM↓ RMSE↓ ERGAS↓
PZRes-Net [17] TIP’21 40.15M 5262G 0.0141s 11059MB 34.963 1.934 35.498 1.935 37.427 1.478 28.234 1.538
MSSJFL [18] HPCC’21 16.33M 175.56G 0.0128s 1349MB 34.966 1.792 33.636 2.245 38.006 1.390 26.893 1.535
Dual-UNet [19] TGRS’21 2.97M 88.65G 0.0127s 2152MB 35.423 1.892 33.183 1.796 38.453 1.548 26.148 1.205
DHIF-Net [20] TCI’22 57.04M 13795.11G 6.005s 29381MB 34.458 1.829 34.769 2.613 39.146 1.239 25.309 1.113
FusFormer [21] TGRS’22 0.18M 11.74G 0.0158s 5964MB 34.217 2.012 35.687 1.996 38.637 1.678 28.674 1.204
HyperTransformer [22] CVPR’22 142.83M 343.96G 0.0252s 8104MB 28.692 3.664 62.231 4.774 32.954 2.568 41.256 3.834
HyperRefiner [23] TJDE’23 19.32M 94.37G 0.0237s 7542MB 33.298 2.129 38.769 2.086 37.654 1.590 29.629 1.403
QRCODE [24] TGRS’24 41.88M 2231.19G 0.2452s 15028MB 35.361 1.623 32.711 2.027 38.948 1.148 24.617 1.429
HyFusion (Ours) 25.53M 641.49G 0.0294s 2572MB 37.421 1.256 27.425 1.172 40.197 1.017 21.455 1.076
II-C Enhanced Reception Field Block
Given the inherent data scarcity in hyperspectral imaging due to hardware limitations and acquisition costs, maximizing the utility of limited available data becomes crucial. The proposed ERFB is specifically designed to address this challenge along with two fundamental issues in hyperspectral image fusion: capturing comprehensive spatial-spectral relationships through long-range dependencies and maximizing feature utilization efficiency. By effectively reusing features and enhancing the receptive field, ERFB enables the network to learn robust representations even with limited training data.
Our design philosophy centers on a dual-enhancement strategy that synergistically combines dense connections for thorough feature reuse with an ISTL for enhanced receptive field and long -range dependency modeling. The dense connection structure ensures that each layer can access and build upon all previous feature representations, effectively preventing information loss and vanishing gradients that commonly occur in deep networks. This is particularly crucial for hyperspectral image fusion, where preserving fine spectral details while maximizing the utility of limited data is essential.
At the core of our ERFB is the ISTL, which consists of two sequential processing paths: W-MSA and SW-MSA. Each path contains LayerNorm, attention (W-MSA or SW-MSA), LayerNorm, and MLP blocks with learnable parameters $\beta_{1}$ and $\beta_{2}$ respectively to adaptively control feature contributions:
$\text{ISTL}(\mathbf{X})=\beta_{1}\cdot\text{Path}_{\text{W-MSA}}(\mathbf{X})+\beta_{2}\cdot\text{Path}_{\text{SW-MSA}}(\mathbf{X})\vspace{-0.1cm}$ (6)
The dense connection pattern in ERFB progressively accumulates and reuses features for reception field enhancement:
$\mathbf{Z}_{j}=\text{ISTL}(\text{Conv}_{3\times 3}(\text{Cat}(\mathbf{Z}_{0},...,\mathbf{Z}_{j-1}))),j=1,2,3,4\vspace{-0.1cm}$ (7)
where $\text{Cat}(\cdot)$ represents the channel-wise concatenation, and $\text{Conv}_{3\times 3}$ fuses the concatenated features with LeakyReLU activation with $0.2$ negative slope. The final stage processes the accumulated features:
$\mathbf{Z}_{5}=\text{Conv}_{3\times 3}(\text{ISTL}(\mathbf{Z}_{4}))\vspace{-0.1cm}$ (8)
With the efficient information flow via dense connectivity within ERFB, and adaptive feature weighting through ISTL, the long-range dependency modeling capability of the proposed HyFusion can be improved, as shown in Figure 1. In the context of limited hyperspectral training data, this architecture proves particularly valuable as it maximizes the utility of available information through feature reuse while maintaining the ability to capture complex spatial-spectral relationships. When incorporated into both SpeNet and SpaNet branches of the DCN architecture, this approach provides a informative representations for high-quality hyperspectral image fusion while maintaining computational and data efficiency.
II-D Task-Specific Optimization
Recognizing the paramount importance of spectral index in HSI, we introduce a Spectral Angle Mapper (SAM) loss function and Stationary Wavelet Transform (SWT) loss to guide the network:
$\ell_{\text{SAM}}=\frac{1}{N}\sum_{n=1}^{N}\cos^{-1}\left(\frac{\mathbf{Y}_{n}^{T}\mathbf{Y}^{*}_{n}}{\|\mathbf{Y}_{n}\|_{2}\cdot\|\mathbf{Y}_{n}^{*}\|_{2}}\right)\vspace{-0.1cm}$ (9)
$\ell_{\text{SWT}}(\mathbf{Y},\mathbf{Y}^{*})=\mathbb{E}\bigl{[}\sum_{j}\lambda_{j}\big{\|}\text{SWT}(\mathbf{Y})_{j}-\text{SWT}(\mathbf{Y}^{*})_{j}\big{\|}_{1}\bigr{]}\vspace{-0.1cm}$ (10)
Finally, the total loss with L1 loss can be written as:
$\ell_{\text{Total}}(\mathbf{Y},\mathbf{Y}^{*})=\ell_{\text{1}}(\mathbf{Y},\mathbf{Y}^{*})+\lambda_{1}\ell_{\text{SAM}}(\mathbf{Y},\mathbf{Y}^{*})+\lambda_{2}\ell_{\text{SWT}}(\mathbf{Y},\mathbf{Y}^{*})\vspace{-0.1cm}$ (11)
III Experimental Results
III-A Experiment Settings
III-A1 Dataset Illustration
The dataset used for performance evaluation in this study was acquired by the Airborne Visible/Infrared Imaging Spectrometer (AVIRIS) sensor [25]. It consisted of 2,078 HR-HSI images, which were randomly partitioned into training, validation, and testing sets for performance evaluation. The training set contained 1,678 images, while the validation and testing sets contained 200 images for each. The spatial and spectral resolutions of the HR-MSI and LR-HSI were $256\times 256\times M_{m}$ and $64\times 64\times 172$, respectively, where $M_{m}$ is either 4 or 6 in our experiments.
III-A2 Implementation Details
The proposed method was implemented using the PyTorch deep learning framework. The batch size was set to 4, and the number of training epochs was fixed to 600 for all experiments involving the proposed method. For the peer methods, the number of training epochs was set according to their default values as specified in their respectively original publications. The ADAM [26] optimizer was used for training, with an initial learning rate of 0.0001. The learning rate was adjusted during the training process using the Cosine Annealing learning decay scheduler. The weights of the penalty terms in the loss function, denoted as $\lambda_{1}$ and $\lambda_{2}$, were both set to 0.01. Standard data augmentation, including random cropping and rotation, is adopted in this paper for all of the evaluated methods.
Figure 3: Performance comparison between several HSI fusion models. (a) Evaluation of model performance with varying training data sizes. (b) Validation performance curves during training with 5% training samples.
III-B Quantitative Results
To evaluate the performance of our hyperspectral image fusion method, we compared it with five other supervised hyperspectral image fusion methods: PZRes-Net [17], MSSJFL [18], Dual-UNet [19] and DHIF-Net [20], FusFormer [21], HyperTransformer [22], HyperRefiner [23], QRCODE [24]. The performance was objectively measured using PSNR, SAM, and RMSE, ERGAS. Experiments were conducted with both 4 and 6 MSI bands. Table I present the results of comparing our proposed method with the state-of-the-art methods.
The experimental results demonstrate that the proposed framework outperforms the compared state-of-the-art methods in terms of spectral reconstruction performance.
III-C Data Efficiency Analysis
To evaluate HyFusion’s ability to learn from limited training data, we conducted experiments comparing our approach with state-of-the-art methods FusFormer [21], HyperRefiner [23] and MSSJFL [18] using 4 bands LR-HSI for fusion. As shown in 3(a), we gradually reduced the percentage of training samples from 100% to 5%. HyFusion outperforms other methods by a significant margin. This demonstrates our model’s strong data efficiency, enabled by the feature reuse through dense connections and enhanced learning capability from ISTL.
The convergence analysis in 3(b) further validates HyFusion’s advantages. Our model not only achieves faster convergence but also reaches higher PSNR values throughout the training process. This rapid convergence can be attributed to the effective spatial-spectral feature extraction through our dual-branch architecture and the improved gradient flow from dense connections. The results highlight that HyFusion’s architectural innovations lead to both better performance and more efficient learning from limited hyperspectral training data.
IV Conclusion
In this paper, we have proposed HyFusion, a novel framework that effectively addresses the challenges of hyperspectral image fusion through two key innovations: the Enhanced Reception Field Block (ERFB) that integrates dense connections with an Improved Swin Transformer Layer (ISTL) for comprehensive feature utilization and stable information propagation, and a Dual-Coupled Network (DCN) architecture that enables efficient cross-domain feature extraction and data utilization for boosting data-scared hyperspectral imagery application. Our experimental results demonstrate that HyFusion achieves state-of-the-art performance in HR-MSI/LR-HSI fusion while exhibiting remarkable data efficiency with limited training samples. The framework’s success in balancing reconstruction quality with computational efficiency makes it particularly valuable for practical hyperspectral imaging applications in resource-constrained environments.